What is the first step to solve this using the elimination method:
5x-2y=-15
3x+3y=12

Found 3 solutions by macston, rothauserc, fractalier:
Answer by macston(5194)

.
Multiply (5x-2y=-15) by 3.
.
15x-6y=-45
.

Answer by rothauserc(4718)

First step is to solve 3x + 3y = 12 for x
x = 4 -y
:
then substitute (4 - y) for x in equation 5x -2y = -15 and solve for y

Answer by fractalier(6550)

I would multiply the second equation by 2/3, giving you
2x + 2y = 8